# Lessons from the Trenches on Reproducible Evaluation of Language Models

### v1

@@ 1 Introduction
In this work, we detail our lessons learned that have been especially beneficial to obtaining useful and rigorous findings. By sharing these lessons, we aim to foster a more robust and reliable evaluation ecosystem. Our contributions are threefold:
  1. 1.
    We highlight several commonly-faced challenges in evaluating language models, including the difficulty of assessing the correctness of natural language responses, challenges in benchmark design, and the dependence upon implementation details that are often obscured or unreported (Section 2).
  2. 2.
    We then discuss best practices we’ve identified to improve how to communicate results and improve evaluation rigor in the language modeling community, despite–or to mitigate the impact of—these challenges (Section 3).
  3. 3.
    To enable researchers and engineers to easily utilize the best practices we have identified, we present  lm-eval, our open-source library for reproducible evaluation of language models (Section 4). lm-eval’s framework defines a flexible API for both model implementation and evaluation task implementation, and then handles all of the work to orchestrate evaluations internally, allowing users to evaluate their chosen combinations of models + tasks, dependent on use case. We additionally present a set of case studies in which lm-eval was used to improve the level of rigor of a set of evaluations (Section 5).
@@ 2.1 Evaluating and Scoring Natural Language Abilities
To address the high costs of manual human evaluation, automated metrics are often used. These offer notable advantages in that they are (theoretically) fully reproducible, far easier and cheaper to compute, and can avoid some of the issues faced by human studies (Wei & Jia, 2021; Freitag et al., 2021; Amidei et al., 2020). Automated metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) seek to directly solve the Key Problem by measuring the distance from a generated response to a gold-standard one, such as via counting the n-gram overlap between the two texts. Heuristic-based metrics such as BLEU (and its derivatives) have flaws (Callison-Burch et al., 2006) and present reproducibility challenges (Marie et al., 2021), but can be useful. More recently, model-based metrics have recently gained momentum through evaluation methods that leverage large language models as a grader (Kim et al., 2024; Wang et al., 2024; Liu et al., 2023b), especially as proxies for human preference evaluation (Zheng et al., 2023), but these are known to be flawed (Wang et al., 2023; Huang et al., 2024; Shen et al., 2023; Zeng et al., 2024; Hu et al., 2024; Liu et al., 2023c; Chen et al., 2024) and suffer from similar reproducibility issues as BLEU, ROUGE, and their variants.
@@ 2.2 Benchmark Design and Validity
While validity is an ongoing problem in language model evaluation, we focus on mitigating other concerns first: as we will describe, lm-eval is designed to ensure measurements are consistent across runs and models, regardless of (construct) validity.
@@ 2.3.2 Lack of Agreement About “Apples to Apples”
Likewise, some questions of how to set up controlled experiments are still open–is it ideal to normalize performance and comparisons by the number of parameters? Training FLOPs? Inference cost? Must training data be held equal? How should models which can leverage external resources such as retrieved documents or external tools be compared? These questions are all context-dependent but can impact findings significantly. For example, Wang et al. (2022) explore comparisons across architectures and training objectives, and choose to normalize for FLOPs, thus comparing encoder-decoder models with double the parameters to decoder-only models. Comparing results of models with equivalent training FLOPs, regardless of the allocation of those FLOPs, is commonplace (Hoffmann et al. (2022); Peng et al. (2023); Touvron et al. (2023), inter alia). However, in a more memory-constrained setting, comparing models equi-parameter may be more logical. While this is not inherently problematic, as different application contexts motivate different evaluation criteria, it is common to gloss headline claims as “model X is better than model y” or “model x doesn’t really work as advertised” without paying significant attention to how comparisons were made.
@@ 2.4 Fast-changing Progress and Conventions
Figure 1: A timeline showing the relative release dates of a selection of notable benchmarks used to evaluate LMs, as compared to the release dates of BERT (Devlin et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and ChatGPT, used as approximate stand-ins for shifts in how the community uses and therefore evaluates LMs.
@@
  • There is no “ground-truth” implementation from the original benchmark authors for many of these popular benchmarks “retrofitted” to be used with prompted autoregressive LMs. In the absence of a clear standard, the community’s methodology for evaluating on these benchmarks may be fragmented or undocumented (Clark et al., 2018; Paperno et al., 2016).
@@ 3 Best Practices for Language Model Evaluation
  • Quantitative scores only provide so much information. To understand why a model is scoring so well or so poorly, it is important to do some sort of qualitative error analysis. This can sometimes reveal superficial errors that are easier to correct with post-processing Bawden & Yvon (2023), or more fundamental errors.
Perform statistical significance testing
  • Most works on language modeling do not perform statistical significance testing (Marie et al., 2021). This simple addition can dramatically boost the reliability of claimed results.
@@ 4 The Language Model Evaluation Harness
Informed by these practices we have built lm-eval. Unlike prior work on unified benchmarking libraries (Liang et al., 2023; Srivastava et al., 2022), the Evaluation Library does not seek to solely prescribe what the correct benchmark or evaluation protocols to use are, and allows users to select their desired tasks and use cases.
The role of the lm-eval is to solve the orchestration problem: previously, performing thorough LM evaluations would require painstaking re-implementation of previous tasks (likely to introduce subtle methodological divergences) or the individual installation and usage of extant code for each benchmark. Our goal is to make it easy to allow researchers or library users to simply install one codebase, and run their method plus selected baselines on their desired tasks in a controlled fashion. We hope to make it more effort-intensive to not follow minimum best practices for evaluation rigor.
@@ 4.2 Addressing Challenges and Incorporating Best Practices
lm-eval reports the standard error (SE) of most supported metrics, calculated by either bootstrapping or dividing the sample standard deviation by the root of the sample size.
By reporting these SE calculations prominently in every evaluation run, we make it trivial for practitioners to add simple statistical measures such as confidence intervals to their results. While we believe more rigorous and widespread statistical testing in LM evaluation is still needed, we hope that this will spur the community to report and be more aware of statistical significance concerns by lowering the difficulty of reporting such measures.
@@ 5 Case Studies
Finally, we demonstrate lm-eval’s utility for improving evaluation rigor and understanding via case studies of its successful usage. We provide additional case studies within Appendix B.
5.1 Prompts Massively Affect Results
As mentioned in Section 2.3.1, language models can be very sensitive to the specific prompt or framing of an evaluation task.
Here we detail how lm-eval has been used to study this sensitivity, and can be used to improve confidence in the comparison of scores across models. We focus our attention on two popular language modeling benchmarks: the ARC question answering benchmark (Clark et al., 2018) and MMLU (Hendrycks et al., 2021). However, not all papers evaluate on these tasks in the same way as the original formats.
ARC Easy ARC Challenge MMLU
Cloze MMLU-style Cloze MMLU-style MMLU-style Hybrid
GPT-NeoX-20B 72.7% 26.3% 38.0% 26.6% 24.5% 27.6%
Llama-2-7B 76.4% 56.2% 43.5% 42.8% 41.3% 39.8%
Falcon-7B 74.6% 28.7% 40.2% 25.9% 25.4% 29.1%
Mistral-7B 80.8% 86.3% 50.1% 72.4% 58.6% 48.3%
Mixtral-8x7B 84.3% 92.1% 56.7% 81.3% 67.1% 59.7%
Table 1: Comparison of 0-shot model performance for several pretrained LMs (Black et al., 2022; Touvron et al., 2023; Penedo et al., 2023; Jiang et al., 2023; 2024) on ARC (Easy and Challenge subsets) and MMLU across two commonly used prompt styles, evaluated using lm-eval.
@@ 5.1 Prompts Massively Affect Results
However, if models do not adopt these approaches, or disclose their exact settings, it is impossible to reliably compare stated model performance. In Table 1, we compare evaluation on the Easy and Challenge sets of ARC using the prompt from Brown et al. (2020) (“Cloze”) and using an MMLU-style answer letter with explicit multiple choice options (“MMLU-style”). We additionally compare MMLU scores between the original MMLU prompting style (“MMLU-style”) and an approach we term “Hybrid”, consisting of an MMLU-style prompt but using the answer strings instead of answer letters as the set of continuations over which we can score. As described further in Appendix B.1, this can be done by modifying two lines in lm-eval’s ARC and MMLU config files. Comparing these prompting styles, the degree to which models differ in performance widely varies, as well as which prompt performs better. If certain model creators chose one prompting style and certain other model creators chose the other, and each used the cited numbers from the respective technical reports for comparing their model to other baselines, the comparisons would be nonsensical and not provide information on which model were “truly” performant.
@@ 5.2 Empowering Benchmark Creators and LM Evaluation Research
lm-eval has recently received contributions for a variety of datasets relying on lm-eval for their evaluation and benchmark prototyping/design (Faysse et al., 2024; Son et al., 2024a; b; Kweon et al., 2024; Li et al., 2024). By directly contributing their new evaluation tasks to lm-eval, benchmark authors also get to have full control over the dissemination (Hendrycks & Woodside, 2024) and implementation of their evaluation, making it far easier for the language modeling community to discover and recognize new benchmarking contributions that might otherwise go unrecognized or unadopted (Dehghani et al., 2021). This is the power of orchestration–the goal is to put new evaluation benchmarks in the hands of the community, and put tools for creating benchmarks given an evaluation dataset in the hands of evaluation developers, while smoothing over potential roadblocks we discuss in Section 2. As an additional concrete example, tasks in lm-eval have been used to back not only the popular Open LLM Leaderboard (Beeching et al., 2023), but also the construction of arbitrary novel leaderboards, which have been used to make custom comparisons between models on more specific use cases, such as non-English languages.
@@ A.2 Ranking-Based Multiple Choice QA
While this factor does not impact the validity of the above calculations, we note that this implies one should be very careful with how their prompt will be tokenized, especially in cases where the tokenization of an input and output pair separately may not align with the tokenization the LM most often saw during training. Mitigations can be considered to remedy this issue, such as “token healing” [^5] https://github.com/guidance-ai/guidance/blob/main/notebooks/tutorials/token_healing.ipynb.
@@
We hope that future work can examine the most practical way to remove such tokenization-based concerns and difficulties from evaluation, such as BPE dropout (Provilkov et al., 2020) or other regularization techniques, or other novel tokenization innovations.
@@ Appendix B Case Studies
B.1 Prompt Sensitivity

### v2

@@ 1 Introduction
Unfortunately, transparent and reproducible evaluation of large language models is very challenging. In our research we have frequently struggled to reproduce the results reported in various papers as well as carry out new evaluations ourselves. To address this problem we built the Language Model Evaluation Harness (Gao et al., 2021), a flexible evaluation library that serves as research infrastructure for evaluation. Our goal with lm-eval is to enable researchers to run any benchmark on any model as easily as possible, while also making it easy for creators of new model inference libraries or evaluation benchmarks to connect their work to the broader ecosystem.
Over the past three years, the design of lm-eval has evolved as the needs of the open source community and our understanding of best practices for language model evaluation have evolved. In this paper we detail lessons learned that have been especially beneficial to obtaining useful and rigorous findings. We highlight several commonly-faced challenges in evaluating language models, including the difficulty of assessing the correctness of natural language responses, challenges in benchmark design, and the dependence upon implementation details that are often obscured or unreported (Section 2). We then discuss best practices we’ve identified to improve how to communicate results and improve evaluation rigor in the language modeling community, despite these challenges (Section 3). Finally, we detail how we have used our learnings to inform the design of lm-eval (Section 4).
@@ 2.1 Evaluating and Scoring Natural Language Abilities
To address the high costs of manual human evaluation, automated metrics are often used. These offer notable advantages in that they are (theoretically) fully reproducible, far easier and cheaper to compute, and can avoid some of the issues faced by human studies (Wei & Jia, 2021; Freitag et al., 2021; Amidei et al., 2020). Automated metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) seek to directly solve the Key Problem by measuring the distance from a generated response to a gold-standard one, such as via counting the n-gram overlap between the two texts. Heuristic-based metrics such as BLEU (and its derivatives) have flaws (Callison-Burch et al., 2006) and present reproducibility challenges (Marie et al., 2021), but can be useful. More recently, model-based metrics have recently gained momentum through evaluation methods that leverage large language models as a grader (Kim et al., 2024; Wang et al., 2024; Liu et al., 2023b), especially as proxies for human preference evaluation (Zheng et al., 2023), but these are known to be flawed (Wang et al., 2023; Shen et al., 2023a; Zeng et al., 2024; Hu et al., 2024; Liu et al., 2023c; Chen et al., 2024) and suffer from similar reproducibility issues as BLEU, ROUGE, and their variants.
@@ 2.2 Benchmark Design and Validity
While validity is an ongoing problem in language model evaluation, we focus on mitigating other concerns first: as we will describe, lm-eval is designed to ensure measurements are consistent across runs and models, regardless of (construct) validity. This is due to our goal of building research infrastructure for evaluations. While we as researchers prefer some evaluation benchmarks to others, our goal is to enable researchers to run any evaluation benchmarks on any models.
@@ 2.3.2 Lack of Agreement About “Apples to Apples”
Likewise, some questions of how to set up controlled experiments are still open–is it ideal to normalize performance and comparisons by the number of parameters? Training FLOPs? Inference cost? Must training data be held equal? How should models which can leverage external resources such as retrieved documents or external tools be compared? These questions are all context-dependent but can impact findings significantly. For example, Wang et al. (2022) explore comparisons across architectures and training objectives, and choose to normalize for FLOPs, thus comparing encoder-decoder models with double the parameters to decoder-only models. Comparing results of models with equivalent training FLOPs, regardless of the allocation of those FLOPs, is commonplace (Hoffmann et al. (2022); Peng et al. (2023); Touvron et al. (2023), inter alia). However, in a more memory-constrained setting, comparing models equi-parameter may be more logical. While this is not inherently problematic, as different application contexts motivate different evaluation criteria, it is common to gloss headline claims as referring to the general case without paying significant attention to such caveats.
@@ 2.4 Fast-changing Progress and Conventions
  2. 2.
    There is no “ground-truth” implementation from the original benchmark authors for many of these popular benchmarks “retrofitted” to be used with autoregressive LMs. In the absence of a clear standard, the community’s methodology for evaluating on these benchmarks may be fragmented or undocumented (Clark et al., 2018; Paperno et al., 2016).
@@
Figure 1: A timeline showing the relative release dates of a selection of notable benchmarks used to evaluate LMs, as compared to the release dates of BERT (Devlin et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and ChatGPT, used as approximate stand-ins for shifts in how the community uses and therefore evaluates LMs. Common practice for evaluating autoregressive language models today diverges from the method described in the paper for all listed tasks except MMLU and MATH.
@@ 3 Best Practices for Language Model Evaluation
  • Quantitative scores only provide so much information. To understand why a model is scoring so well or so poorly, it is important to do some sort of qualitative error analysis. This can sometimes reveal superficial errors that are easier to correct with post-processing (Bawden & Yvon, 2023), or more fundamental errors.
Measure and Report Uncertainty
  • Most works on language modeling do not perform statistical significance testing, which can significantly decrease confidence in results (Marie et al., 2021).
@@
  • Even when not retraining models, statistical analysis can bound the expected variation across model training runs (Jordan, 2023).
@@ 4 The Language Model Evaluation Harness
Informed by these practices we have built lm-eval. Unlike subsequent555lm-eval was originally built in 2021 and has been in continuous use at EleutherAI and elsewhere since then, despite not being formally introduced in any papers. work on unified benchmarking libraries (Liang et al., 2023; Srivastava et al., 2022; Barton, 2024), the Evaluation Harness does not seek to solely prescribe what the correct benchmark or evaluation protocols to use are, and allows users to select their desired tasks and use cases.
The role of the lm-eval is to solve the orchestration problem: previously, performing thorough LM evaluations would require painstaking re-implementation of previous tasks (likely to introduce subtle methodological divergences) or individually installing, debugging, and using dozens of small libraries. Our goal is to make it easy to allow researchers or library users to simply install one codebase, and run their method plus selected baselines on their desired tasks in a controlled fashion. At the same time, we strive to design best practices into the functionality of the library itself, so that the default and easy-to-use functionality guides users to follow best practices.
@@ 4.2 Addressing Challenges and Incorporating Best Practices
lm-eval reports the standard error (SE) of most supported metrics [^6] The standard error is always the standard error of the evaluation metric. In most, but not all, cases this is the standard error of the mean, as most evaluation benchmarks report the mean score as their final metric., calculated by either bootstrapping or dividing the sample standard deviation by the root of the sample size.
By reporting these SE calculations prominently in every evaluation run, we make it easy for practitioners to add simple statistical measures such as confidence intervals to their results. While we believe more rigorous and widespread statistical testing in LM evaluation is still needed, we hope that this will spur the community to report and be more aware of statistical significance concerns and lower the difficulty of reporting such measures.
Note that the standard errors described here refer to an estimate of the variance in the observed scores should the evaluation data be recollected from the same distribution (Recht et al., 2019; Zhang et al., 2024). Other forms of variance, such as temporal drift (Longpre et al., 2023), retraining models with different seeds (a paradigm we are currently experimenting with), or variance across prompts (Sanh et al., 2022; Muennighoff et al., 2022) must be calculated differently.
@@ 5 Case Studies
Finally, we demonstrate lm-eval’s utility for improving evaluation rigor and understanding via case studies of its successful usage. Additional case studies can be found in Appendix B.
5.1 Multiprompt Evaluation with the BigScience Workshop
Even when scoring criteria are held identical, the specific prompt of an evalution task can heavily impact results. In collaboration with the BigScience Workshop, the PromptSource (Bach et al., 2022) library was added to a fork of lm-eval to enable easy evaluation across many different prompt templates [^7] https://github.com/bigscience-workshop/lm-evaluation-harness for the first time. Most papers that came out of the BigScience Workshop used this functionality to report distributions of scores across different prompting set-ups (Sanh et al., 2022; Muennighoff et al., 2022; Yong et al., 2023; Workshop et al., 2023). PromptSource, along with other innovations introduced in the BigScience fork, is now supported natively in lm-eval. We have also further extended our functionality to enable people to use the Jinja templating language directly in their configuration files to make it easy to define custom evaluation templates by hand and algorithmically, regardless of whether they use prompts from the PromptSource library.
Figure 4: An example of using lm-eval to study the impact of prompts on model performance, taken from Muennighoff et al. (2022). Dots represent different prompts, as summarized by the histograms. Similar plots can be found in Workshop et al. (2023); Sanh et al. (2022); and others.
While the approach used by BigScience in reporting distributions across prompts has not been widely adopted, the idea that prompts should be considered as part of the evaluation set-up has become widely accepted. We hope that future research will especially continue to focus on collecting realistic prompts (Shen et al., 2023b; Xie et al., 2023; Hofmann et al., 2024) or measuring the extent to which results with particular common set-ups generalize to more realistic use-cases (Lyu et al., 2024), or otherwise investigate prompting as a human-computer interaction problem.
5.2 Difficulties in Comparing Scores Across Evaluation Setups
As mentioned in Section 2.3, scores on evaluation tasks can be substantially affected by the specifics of the evaluation implementation and setup. Here, we provide an example of studying such sensitivity to divergences in evaluation methodology, and how lm-eval can be used to improve confidence in the comparison of scores across models by preventing such divergences. We focus our attention on two popular language modeling benchmarks: the ARC question answering benchmark (Clark et al., 2018) and MMLU (Hendrycks et al., 2021).
ARC Challenge MMLU
Cloze MMLU-style MMLU-style Hybrid
GPT-NeoX-20B $38.0\pm 2.78$% $26.6\pm 2.53$% $24.5\pm 0.71$% $27.6\pm 0.74$%
Llama-2-7B $43.5\pm 2.84$% $42.8\pm 2.83$% $41.3\pm 0.80$% $39.8\pm 0.79$%
Falcon-7B $40.2\pm 2.81$% $25.9\pm 2.51$% $25.4\pm 0.72$% $29.1\pm 0.75$%
Mistral-7B $50.1\pm 2.86$% $72.4\pm 2.56$% $58.6\pm 0.77$% $48.3\pm 0.80$%
Mixtral-8x7B $56.7\pm 2.84$% $81.3\pm 2.23$% $67.1\pm 0.72$% $59.7\pm 0.77$%
Table 1: Comparison of 0-shot model performance for several pretrained LMs (Black et al., 2022; Touvron et al., 2023; Penedo et al., 2023; Jiang et al., 2023; 2024) on ARC (Challenge subset) and MMLU across two commonly used prompt styles, evaluated using lm-eval. Unnormalized accuracy is reported, alongside 95% confidence intervals via SE of the mean.
@@ 5.2 Difficulties in Comparing Scores Across Evaluation Setups
However, if models do not adopt these approaches, or disclose their exact settings, it is impossible to reliably compare stated model performance. In Table 1, we compare evaluation on the Challenge subset of ARC using the prompt from Brown et al. (2020) (“Cloze”) and using an MMLU-style answer letter with explicit multiple choice options (“MMLU-style”) [^8] While they do not report their evaluation set-up, this latter style appears to produce scores consistent with those reported in Jiang et al. (2024) for 25-shot ARC Challenge. We were unable to find a way to reproduce their scores using the standard cloze-style evaluation.. We additionally compare MMLU scores between the original MMLU prompting style (“MMLU-style”) and an approach we term “Hybrid”, consisting of an MMLU-style prompt but using the answer strings instead of answer letters as the set of continuations over which we can score. As described further in Appendix B.2, this can be done by modifying two lines in lm-eval’s ARC and MMLU config files.
Comparing these prompting styles, the degree to which models differ in performance widely varies, as well as which prompt performs better. If certain model creators chose one prompting and scoring style and certain other model creators chose the other, and each used the cited numbers from the respective technical reports for comparing their model to other baselines, the comparisons would be nonsensical and not provide information on which model were “truly” performant. Additionally, better statistical reporting such as the use of confidence intervals which we report, does not resolve these issues–while it gives a sense of how reliable a given measurement is (for example, MMLU has a smaller confidence interval due to the use of a larger amount of samples), it cannot tell us how much model performance will vary across different measurement settings and cannot indicate a comparison should not be made.
@@ 5.3 Empowering Benchmark Creators and LM Evaluation Research
lm-eval has recently received contributions for a variety of datasets relying on lm-eval for their evaluation and benchmark prototyping and design (Faysse et al., 2024; Son et al., 2024a; b; Kweon et al., 2024; Li et al., 2024). By directly contributing their new evaluation tasks to lm-eval, benchmark authors also get to have full control over the dissemination (Hendrycks & Woodside, 2024) and implementation of their evaluation, making it far easier for the language modeling community to discover and recognize new benchmarking contributions that might otherwise go unrecognized or unadopted (Dehghani et al., 2021). This is the power of orchestration–the goal is to put new evaluation benchmarks in the hands of the community, and put tools for creating benchmarks given an evaluation dataset in the hands of evaluation developers, while smoothing over potential roadblocks we discuss in Section 2. As an additional concrete example, tasks in lm-eval have been used to back not only the popular Open LLM Leaderboard (Beeching et al., 2023), but also the construction of arbitrary novel leaderboards, which have been used to make custom comparisons between models on more specific use cases, such as non-English languages.
@@ A.2 Ranking-Based Multiple Choice QA
While this factor does not impact the validity of the above calculations, we note that this implies one should be very careful with how their prompt will be tokenized, especially in cases where the tokenization of an input and output pair separately may not align with the tokenization the LM most often saw during training. There are some recently proposed mitigations to remedy this issue, such as “token healing” [^9] https://github.com/guidance-ai/guidance/blob/main/notebooks/tutorials/token_healing.ipynb,
@@
We hope that future work can examine the most practical way to remove such tokenization-based concerns and difficulties from evaluation, such as BPE dropout (Provilkov et al., 2020), other regularization techniques, or other novel tokenization innovations.
@@ Appendix B Case Studies
B.1 Case Study: Benchmarking Novel LM Architectures
The selection of tasks we check are the following:
Wikitext-103 (“Wiki”)
(Merity et al., 2016): Wikitext-103 is a 103 million word language modeling dataset sourced from Wikipedia by Merity et al. (2016) to serve as a language modeling benchmark. It contains a training, validation, and test split, with a typical setup being to train a (small) model from scratch on the dataset and evaluate its test set perplexity (PPL).
Long Range Arena (“LRA”)
(Tay et al., 2021): LRA is a sequence modeling dataset consisting of a suite of various tasks meant to test the long-range modeling abilities of models. While solving the more challenging longer-context tasks in LRA drove earlier work on long-context sequence models (Gu et al., 2022), subsequent work has shown that LRA may not correlate with desired downstream tasks for pretrained models (Alam et al., 2024).
PIQA (“PQ”)
(Bisk et al., 2020): PIQA is a question-answering dataset meant to evaluate physical common-sense reasoning. It is typically evaluated using loglikelihood-based multiple choice and normalized (acc_norm) or unnormalized (acc) accuracy is reported.
OpenBookQA (“OB”)
(Mihaylov et al., 2018): OpenBookQA is a question-answering dataset meant to evaluate the combination of common knowledge with open-book exam questions. It is also typically evaluated using loglikelihood-based multiple choice and normalized (acc_norm) or unnormalized (acc) accuracy is reported.
WinoGrande (“WG”)
(Sakaguchi et al., 2019): WinoGrande is a dataset consisting of Winograd Schema Challenge-like minimal sentence pairs with one word flipped. Language models are typically evaluated on this dataset by comparing the probability of correctly completing the end of the sentence given the correct or incorrect context (the sentence up to and including the flipped word), and reporting accuracy (acc) (Radford et al., 2019; Brown et al., 2020).
HellaSwag (“HS”)
(Zellers et al., 2019): HellaSwag is an adversarially created dataset meant to test “commonsense natural language inference” mined from WikiHow. Models are typically evaluated by choosing the most likely to be generated completion text from a correct option and (nonsensical) set of incorrect answer options (acc, acc_norm).
AI2 Reasoning Challenge (“ARC”)
(Clark et al., 2018): ARC is a challenging question answering dataset consisting of an Easy and Challenge subset. Questions are sourced from standardized tests on natural sciences. lm-eval follows Brown et al. (2020) in using a “cloze” style loglikelihood-based evaluation and reports acc, acc_norm over the set of answer strings.
LAMBADA (“LMB”)
(Paperno et al., 2016): The LAMBADA dataset is a word prediction benchmark consisting of short passages from Book Corpus (Zhu et al., 2015) books, with a language model required to predict the final word. Radford et al. (2019) introduce a cleaned and detokenized variant of LAMBADA  [^13] For more information, see here and here., often denoted as “Lambada (OpenAI).”
SuperGLUE (“SGLUE”)
(Wang et al., 2019a): SuperGLUE is a benchmark containing a collection of NLU tasks (BoolQ (Clark et al., 2019), CB (De Marneffe et al., 2019), COPA (Roemmele et al., 2011), MultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), RTE (Wang et al., 2019b; Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), WiC (Pilehvar & Camacho-Collados, 2018), WSC (Levesque et al., 2012)). lm-eval implements SuperGLUE as a loglikelihood-based multiple choice classification task based on (Brown et al., 2020).
Arch. Wiki LRA PQ OB WG HS ARC LMB SGLUE
H3 (Fu et al., 2023)
Hyena (Poli et al., 2023)
Pythia (Biderman et al., 2023)
RWKV-4 (Peng et al., 2023)
RetNet (Sun et al., 2023)
HGRN1 (Qin et al., 2024b)
Mamba (Gu & Dao, 2023)
GLA (Yang et al., 2024)
Based (Arora et al., 2024)
Griffin,Hawk (De et al., 2024)
RWKV-5,6 (Peng et al., 2024)
HGRN2 (Qin et al., 2024a)
B.2 Comparisons Across Evaluation Settings